OpenStudy (anonymous):

Solve for y, then find the value of y when given x = –2. 6x = 7 – 4y

OpenStudy (vishweshshrimali5):

-12 = 7 - 4y => y = 19/4

OpenStudy (anonymous):

6x-7=-4y y=(-6x+7)/4 y=19/4

OpenStudy (anonymous):

-19=14y y=19/4